# Effects of bumblebee gravity on neutrino motion

### v1

@@ I Introduction
The search for generalized black hole solutions beyond the standard (A)dS–Schwarzschild spacetime has led to a range of theoretical extensions, including scenarios in which the conventional vacuum assumptions are relaxed, resulting in modified geometric and dynamical properties of spacetime 20 . A notable refinement of this idea arises in the context of bumblebee gravity, where a vector field spontaneously acquires a non–zero expectation value in the temporal direction. This mechanism triggers the breaking of local Lorentz invariance and yields new classes of black hole configurations that deviate from those predicted by general relativity 24 ; 21 ; 23 ; 22 . These alternative backgrounds provide a framework to examine gravitational phenomena under symmetry-violating conditions.
Neutrinos stand out from other fundamental particles due to their unconventional quantum characteristics, which have attracted significant attention in particle physics analysis neu42 ; neu43 ; neu44 . One of their most notable features is that the states responsible for weak interactions—referred to as flavor states—do not coincide with the states that possess definite mass. This misalignment causes neutrinos to evolve as a mixture of mass eigenstates, leading to an oscillatory pattern in their flavor composition as they propagate through space neu40 ; neu39 ; neu41 . The resulting flavor change, known as neutrino oscillation, arises from quantum interference between these mass components.
@@
Gravitational fields can markedly influence neutrino oscillation dynamics, introducing effects absent in flat spacetime scenarios. While in flat geometry the flavor transitions depend solely on differences between mass–squared values, curved spacetime geometries modify the oscillation phase in a way that can reveal sensitivity to the absolute mass scale of neutrinos Shi:2024flw ; AraujoFilho:2025rzh ; Shi:2025rfq . These curvature–induced corrections alter the conventional phase relations and become especially significant for highly energetic neutrinos emitted from astrophysical or cosmological sources Shi:2023kid . As neutrinos propagate through regions with nontrivial geometry, the accumulated phase shift encodes information about both their internal parameters and the structure of the spacetime they traverse. Observational discrepancies between expected flavor distributions and those predicted by standard oscillation models can thus indicate gravitational influences. In this manner, this approach provides a pathway to explore the mass spectrum of neutrinos while simultaneously probing the nature of the gravitational environments encountered during their journey neu53 ; neu46 ; neu49 ; neu48 ; neu47 ; neu50 ; neu51 ; Shi:2023hbw ; neu52 .
Neutrino flavor dynamics are strongly shaped by the geometry of the spacetime they traverse. When described in geometric terms, the phase acquired along their trajectories reflects the influence of the gravitational background neu55 ; neu54 , as one should naturally expect. In highly curved environments—such as near compact massive bodies—gravitational lensing can distort neutrino paths, causing them to converge or intersect. This bending alters the interference between different mass eigenstates, leading to modifications in the oscillation behavior and affecting the resulting flavor transition probabilities Shi:2025rfq ; neu53 ; Shi:2024flw .
@@ II The bumblebee solution in the metric formalism
$\mathrm{d}s^{2}=$ $-\left(1-\dfrac{2M}{r}\right)\mathrm{d}t^{2}+(1-\ell)\left(1-\dfrac{2M}{r}\right)^{-1}\mathrm{d}r^{2}$ (1)
@@
×$\times R^{-\frac{3}{2}}\int_{1}^{\infty}(x-1)^{4}\left(x^{2}+4x+5\right)\sqrt{\dfrac{\mathfrak{g}_{rr}(yR)}{-\mathfrak{g}_{tt}^{9}(yR)}}y^{2}\mathrm{d}y.$ (19)
@@
Figure 2: The radial profile of the differential energy deposition rate, $\mathrm{d}\dot{Q}/\mathrm{d}r$, is analyzed for different values of the compactness ratio $M/R$. In the Newtonian regime—corresponding to the case where $M=0$—the relation reduces to a simpler form, resulting in $\mathrm{d}\dot{Q}/\mathrm{d}r=1$ precisely at the boundary $r=R$.
@@
To investigate the impact of Lorentz symmetry breaking on neutrino–antineutrino annihilation, we reconstruct the formalism for energy deposition using a Schwarzschild–like solution modified within the context of Bumblebee gravity. This approach allows us to analyze deviations from general relativity induced by the parameter $\ell$. Fig. 1 presents the results for the energy conversion efficiency ratio, given by Eq. (20), for a range of $\ell$ values. The comparison between the standard relativistic case and the Lorentz–violating extensions reveals an unexpected enhancement in the energy deposition rate across all curves, despite the parameter $\ell$ varying from 00 to $0.3$. Notably, these enhancements preserve similar radial behavior when compared to general relativity. This result contrasts with the suppression patterns found in gravitational modifications arising from the Kalb–Ramond gravity recently addressed in the literature Shi:2025rfq .
@@
A comparison between the two gravitational models uncovers contrasting mechanisms that shape energy deposition behavior. In other words, in the Bumblebee scenario, the Lorentz–violating parameter $\ell$ introduces modifications through the coupling of the background vector field with matter content, generating effects that resemble repulsive corrections. These alterations partially mitigate energy loss in regions dominated by intense gravitational fields. In contrast, the Kalb–Ramond model amplifies the suppression of energy conversion by deepening the gravitational potential via tensor field contributions, thereby reducing the likelihood of efficient neutrino–antineutrino interactions. The results point to the possibility that Lorentz–violating effects in the Bumblebee model could open up a new pathway for energy production in extremely dense astrophysical systems, like the remnants left behind after neutron star mergers. This mechanism may help explain sudden spikes in brightness often seen in short–duration gamma–ray bursts.
@@ IV Neutrino phase evolution and conversion probability
To investigate the trajectory of neutrinos linked to the $k$–th eigenmode in a spacetime exhibiting spherical symmetry, one may adopt the Lagrangian methodology presented in Ref. neu18 . This framework, applied to the metric structure introduced in Eq. (1), enables a systematic derivation of the governing equations for neutrino motion by exploiting variational principles adapted to curved backgrounds as follows
@@
By confining the motion to the equatorial plane, where $\theta=\pi/2$, the resulting dynamics involve only a limited set of non–zero momentum components. To obtain these, one starts by introducing the canonical momentum $p_{\mu}$ through the relation $p_{\mu}=\partial\mathcal{L}/\partial(\mathrm{d}x^{\mu}/\mathrm{d}\tau)$, where $\mathcal{L}$ denotes the Lagrangian and $\tau$ represents the proper time parameter. The quantity $m_{k}$ designates the mass associated with the $k$–th neutrino eigenstate. Detailed forms of the relevant momentum components that survive under this symmetry reduction are provided in Refs. neu60 ; Shi:2024flw
@@
Notice that, as one should expect, the relativistic dynamics of a neutrino belonging to the $k$–th mass eigenstate are governed by the mass–shell relation, which establishes the constraint $\mathfrak{g}^{\mu\nu}p_{\mu}p_{\nu}=-m_{k}^{2}$. This condition guarantees compatibility between the particle’s motion and the underlying spacetime geometry, ensuring that its trajectory remains consistent with general relativistic principles neu54 ; neu55
@@
|$\ket{\nu_{\alpha}}=\sum U_{\alpha i}^{*}\ket{\nu_{i}}.$ (28)
The evolution of neutrinos through spacetime is most effectively characterized using mass eigenstates, labeled as |$\ket{\nu_{i}}$, each following a unique trajectory influenced by its mass. Despite this, neutrinos are produced and measured in flavor states—$\nu_{e}$, $\nu_{\mu}$, and $\nu_{\tau}$—indexed by $\alpha=e,\mu,\tau$, due to the nature of weak interactions. The connection between these physical flavor states and the underlying mass basis is governed by a unitary transformation involving a $3\times 3$ mixing matrix $U$, as outlined in Ref. neu41 . To model their propagation, one typically associates spacetime coordinates $\left(t_{S},\bm{x}_{S}\right)$ and $\left(t_{D},\bm{x}_{D}\right)$ with the emission and detection points, respectively. Each mass eigenstate wave packet then evolves along the spacetime trajectory linking these two locations
|$\ket{\nu_{i}\left(t_{D},\bm{x}_{D}\right)}=\exp({-\mathbbm{i}\tilde{\Phi}_{i}})\ket{\nu_{i}\left(t_{S},\bm{x}_{S}\right)}.$ (29)
@@
=±$=\pm\frac{m_{k}^{2}}{2E_{0}}\int_{r_{S}}^{r_{D}}\sqrt{-\mathfrak{g}_{tt}\mathfrak{g}_{rr}}\left(1-\dfrac{b^{2}|\mathfrak{g}_{tt}|}{\mathfrak{g}_{\varphi\varphi}}\right)^{-\frac{1}{2}}\mathrm{d}r.$ (33)
@@
$\simeq\left(1-\dfrac{\ell}{2}\right)\left[\dfrac{r}{\sqrt{r^{2}-b^{2}}}-\dfrac{b^{2}M}{\left(r^{2}-b^{2}\right)^{\frac{3}{2}}}\right].$ (34)
@@
$\tilde{\Phi}_{k}$ $=\dfrac{m_{k}^{2}}{2E_{0}}\left(1-\dfrac{\ell}{2}\right)\Biggl{[}\sqrt{r_{D}^{2}-b^{2}}-\sqrt{r_{S}^{2}-b^{2}}$
@@
Within this approach, the average energy of relativistic neutrinos originating from the source is given by $E_{0}=\sqrt{E_{k}^{2}-m_{k}^{2}}$, where $E_{k}$ denotes the energy and $m_{k}$ the mass of the $k$–th mass eigenstate. The quantity $b$, known as the impact parameter, is addressed in Ref. neu18 . As the neutrinos move through the curved background, they approach the central object down to a minimum radial coordinate $r=r_{0}$. In the regime of a weak gravitational field, the equation determining this point of closest approach can be solved analytically using appropriate approximations
@@
$\simeq\frac{{m}_{k}^{2}}{2E_{0}}\left(1-\dfrac{\ell}{2}\right)\Biggl{[}\sqrt{r_{D}^{2}-r_{0}^{2}}+\sqrt{r_{S}^{2}-r_{0}^{2}}$
@@
$\tilde{\Phi}_{k}$ $\simeq\frac{{m}_{k}^{2}}{2E_{0}}\left(1-\dfrac{\ell}{2}\right)\Biggl{[}\sqrt{r_{D}^{2}-b^{2}}+\sqrt{r_{S}^{2}-b^{2}}$
@@
$\tilde{\Phi}_{k}=\dfrac{m_{k}^{2}}{2E_{0}}\left(1-\dfrac{\ell}{2}\right)(r_{D}+r_{S})\left(1-\dfrac{b^{2}}{2r_{D}r_{S}}+\dfrac{2M}{r_{D}+r_{S}}\right).$ (40)
As the Lorentz–violating parameter $\ell$ increases, one observes a clear reduction in the phase acquired by neutrinos during their journey. For the purposes of this evaluation, the parameters are set to $E_{0}=10\,\mathrm{MeV}$, $r_{D}=10\,\mathrm{km}$, and $r_{S}=10^{5}r_{D}$.
@@
$A_{pq}$ $=\frac{r_{S}+r_{D}}{2E_{0}}\left(1-\dfrac{\ell}{2}\right)\left[1+\dfrac{2M}{r_{D}+r_{S}}-\dfrac{\sum b_{pq}^{2}}{4r_{D}r_{S}}\right],$ (44)
$B_{ij}$ $=-\frac{\sum m_{ij}^{2}}{8E_{0}}\left(1-\dfrac{\ell}{2}\right)\left(\frac{1}{r_{D}}+\frac{1}{r_{S}}\right),$ (45)
@@
To distinguish between the phases associated with various neutrino paths, a superscript notation is adopted—$\tilde{\Phi}_{i}^{p}$—where each label $p$ denotes a distinct trajectory characterized by its corresponding impact parameter $b_{p}$. In our case, the resulting phase differences responsible for oscillation phenomena depend on the masses $m_{i}$ of the neutrino eigenstates, the mass–squared differences $\Delta m_{ij}^{2}$, and the geometric features of the spacetime. Notably, in the limit $\ell\to 0$, the expression for the phase shift converges to the standard result reported in Ref. neu53 .
@@ V Gravitational deflection of neutrino trajectories
Now, our analysis turns to the effect of gravitational lensing on neutrino flavor conversion, emphasizing the influence due to $\ell$. Within this lensing configuration, considering a two–flavor neutrino system, the probability of transition from $\nu_{\alpha}$ to $\nu_{\beta}$ is determined under the weak–field approximation, taking into account the spatial arrangement of the emission point, the detector and lensing region neu53 ; neu54 ; neu55 ; Shi:2024flw ; neu65
@@
×$\quad\times\left[\exp\left(-\mathbbm{i}\Delta m_{ij}^{2}A_{11}\right)+\exp\left(-\mathbbm{i}\Delta m_{ij}^{2}A_{22}\right)\right]$
@@
×$\quad\times\exp\left(-\mathbbm{i}\Delta b_{12}^{2}B_{ij}\right)\exp\left(-\mathbbm{i}\Delta m_{ij}^{2}A_{12}\right)$
@@
×$\quad\times\exp\left(\mathbbm{i}\Delta b_{21}^{2}B_{ij}\right)\exp\left(-\mathbbm{i}\Delta m_{ij}^{2}A_{21}\right)\biggr{\}}.$ (51)
@@
×$\quad\times\biggl{[}\sin^{2}\left(\dfrac{1}{2}\Delta m_{12}^{2}A_{11}\right)+\sin^{2}\left(\dfrac{1}{2}\Delta m_{12}^{2}A_{22}\right)$

### v3

@@ I Introduction
The search for generalized black hole solutions beyond the standard (A)dS–Schwarzschild spacetime has led to a range of theoretical extensions, including scenarios in which the conventional vacuum assumptions are relaxed, resulting in modified geometric and dynamical properties of spacetime 20 . A notable refinement of this idea arises in the context of bumblebee gravity, where a vector field spontaneously acquires a non–zero expectation value in the temporal direction. This mechanism triggers the breaking of local Lorentz invariance and yields new possibilities beyond general relativity 24 ; 21 ; 23 ; lambiase2023probing ; 22 ; ovgun2019exact ; Magalhaes:2025nql ; marques2023braneworlds ; engels2016bumblebee ; kanzi2021greybody ; mangut2023probing ; mangut2025lorentz ; Magalhaes:2025lti .
Although quark mixing and meson oscillations are also well established, neutrinos are unique because their tiny masses and flavor–mass misalignment allow coherent oscillations to persist over macroscopic and even astrophysical distances Pontecorvo1 ; Pontecorvo2 ; maki1962remarks . Owing to this property, they have become especially sensitive probes of new physics and gravitational effects neu42 ; neu43 ; neu44 . One of their most notable features is that the states responsible for weak interactions—referred to as flavor states—do not coincide with the states that possess definite mass. This misalignment causes neutrinos to evolve as a mixture of mass eigenstates, leading to an oscillatory pattern in their flavor composition as they propagate through space neu40 ; neu39 ; neu41 . The resulting flavor change, known as neutrino oscillation, arises from quantum interference between these mass components.
@@
Gravitational fields can markedly influence neutrino oscillation dynamics, introducing effects absent in flat spacetime scenarios. While in flat geometry the flavor transitions depend solely on differences between mass–squared values, curved spacetime geometries modify the oscillation phase in a way that can reveal sensitivity to the absolute mass scale of neutrinos Chakrabarty:2023kld ; Alloqulov:2024sns ; Shi:2024flw ; AraujoFilho:2025rzh ; Shi:2025rfq . These curvature–induced corrections alter the conventional phase relations and become especially significant for highly energetic neutrinos emitted from astrophysical or cosmological sources Shi:2023kid . As neutrinos propagate through regions with nontrivial geometry, the accumulated phase shift encodes information about both their internal parameters and the structure of the spacetime they traverse. Observational discrepancies between expected flavor distributions and those predicted by standard oscillation models can thus indicate gravitational influences. In this manner, this approach provides a pathway to explore the mass spectrum of neutrinos while simultaneously probing the nature of the gravitational environments encountered during their journey neu53 ; neu46 ; neu49 ; neu48 ; neu47 ; neu50 ; neu51 ; Shi:2023hbw ; neu52 .
Neutrino flavor dynamics are strongly shaped by the geometry of the spacetime they traverse. When described in geometric terms, the phase acquired along their trajectories reflects the influence of the gravitational background neu54 , as one should naturally expect. In highly curved environments—such as near compact massive bodies—gravitational lensing can distort neutrino paths, causing them to converge or intersect. This bending alters the interference between different mass eigenstates, leading to modifications in the oscillation behavior and affecting the resulting flavor transition probabilities Shi:2025rfq ; neu53 ; Shi:2024flw .
@@ II The bumblebee solution in the metric formalism
$\mathrm{d}s^{2}=$ $-\left(1-\dfrac{2M}{r}\right)\mathrm{d}t^{2}+(1+\ell)\left(1-\dfrac{2M}{r}\right)^{-1}\mathrm{d}r^{2}$ (1)
@@
$\times R^{-\frac{3}{2}}\int_{1}^{\infty}(x-1)^{4}\left(x^{2}+4x+5\right)\sqrt{\dfrac{\mathfrak{g}_{rr}(yR)}{-\mathfrak{g}_{tt}^{9}(yR)}}y^{2}\mathrm{d}y.$ (19)
@@
Figure 2: The radial profile of the differential energy deposition rate, $\mathrm{d}\dot{Q}/\mathrm{d}r$, is analyzed for different values of the compactness ratio $M/R$. In the Newtonian regime—corresponding to the case where M=0M=0—the relation reduces to a simpler form, resulting in $\mathrm{d}\dot{Q}/\mathrm{d}r=1$ precisely at the boundary $r=R$.
@@
Clearly, the integrand in Eq. (20) is smooth and non-oscillatory, and the integration interval $1,\infty]$ is free of discontinuities. The endpoints $y\to 1$ and $y\to\infty$ have limit values of $5R^{5}\sqrt{1+\ell}/(R-2)^{5}$ and 0, respectively. The Simpson’s integration technique may be used to evaluate such an integrand, and the Python utility scipy.integrate can be used to acquire the numerical integration result. We have set the relative and absolute error tolerances to $1.49\times 10^{-8}$. It is predicted that the absolute error of the outcomes throughout the numerical computing process ranges from $3.8\times 10^{-11}$ to $4.8\times 10^{-11}$. To investigate the impact of Lorentz symmetry breaking on neutrino–antineutrino annihilation, we reconstruct the formalism for energy deposition using a Schwarzschild–like solution modified within the context of Bumblebee gravity. This approach allows us to analyze deviations from general relativity induced by the parameter $\ell$. Fig. 1 presents the results for the energy conversion efficiency ratio, given by Eq. (20), for a range of $\ell$ values. The comparison between the standard relativistic case and the Lorentz–violating extensions reveals an unexpected enhancement in the energy deposition rate across all curves, despite the parameter $\ell$ varying from 0 to $0.3$. Notably, these enhancements preserve similar radial behavior when compared to general relativity. This result contrasts with the suppression patterns found in gravitational modifications arising from the Kalb–Ramond gravity recently addressed in the literature Shi:2025rfq .
@@
A comparison between the two gravitational models uncovers contrasting mechanisms that shape energy deposition behavior. In other words, in the Bumblebee scenario, the Lorentz–violating parameter $\ell$ introduces modifications through the coupling of the background vector field. These alterations partially mitigate energy loss in regions dominated by intense gravitational fields. In contrast, the Kalb–Ramond model amplifies the suppression of energy conversion by deepening the gravitational potential via tensor field contributions, thereby reducing the likelihood of efficient neutrino–antineutrino interactions. The results point to the possibility that Lorentz–violating effects in the Bumblebee model could open up a new pathway for energy production in extremely dense astrophysical systems, like the remnants left behind after neutron star mergers. This mechanism may help explain sudden spikes in brightness often seen in short–duration gamma–ray bursts Poddar:2022svz .
@@ IV Neutrino phase evolution and conversion probability
To investigate the trajectory of neutrinos linked to the kk–th eigenmode in a spacetime exhibiting spherical symmetry, one may adopt the Lagrangian methodology presented in Ref. neu18 . This framework, applied to the metric structure introduced in Eq. (1), enables a systematic derivation of the governing equations for neutrino motion by exploiting variational principles adapted to curved backgrounds as follows
@@
By confining the motion to the equatorial plane, where $\theta=\pi/2$, the resulting dynamics involve only a limited set of non–zero momentum components. To obtain these, one starts by introducing the canonical momentum $p_{\mu}$ through the relation $p_{\mu}=\partial\mathcal{L}/\partial(\mathrm{d}x^{\mu}/\mathrm{d}\tau)$, where $\mathcal{L}$ denotes the Lagrangian and $\tau$ represents the proper time parameter. The quantity $m_{k}$ designates the mass associated with the kk–th neutrino eigenstate. Detailed forms of the relevant momentum components that survive under this symmetry reduction are provided in Refs. neu60 ; Shi:2024flw
@@
Notice that, as one should expect, the relativistic dynamics of a neutrino belonging to the kk–th mass eigenstate are governed by the mass–shell relation, which establishes the constraint $\mathfrak{g}^{\mu\nu}p_{\mu}p_{\nu}=-m_{k}^{2}$. This condition guarantees compatibility between the particle’s motion and the underlying spacetime geometry, ensuring that its trajectory remains consistent with general relativistic principles neu54
@@
[…]
@@
$=\pm\frac{m_{k}^{2}}{2E_{0}}\int_{r_{S}}^{r_{D}}\sqrt{-\mathfrak{g}_{tt}\mathfrak{g}_{rr}}\left(1-\dfrac{b^{2}|\mathfrak{g}_{tt}|}{\mathfrak{g}_{\varphi\varphi}}\right)^{-\frac{1}{2}}\mathrm{d}r.$ (33)
@@
$\simeq\left(1+\dfrac{\ell}{2}\right)\left[\dfrac{r}{\sqrt{r^{2}-b^{2}}}-\dfrac{b^{2}M}{\left(r^{2}-b^{2}\right)^{\frac{3}{2}}}\right].$ (34)
@@
$\tilde{\Phi}_{k}$ $=\dfrac{m_{k}^{2}}{2E_{0}}\left(1+\dfrac{\ell}{2}\right)\Biggl[\sqrt{r_{D}^{2}-b^{2}}-\sqrt{r_{S}^{2}-b^{2}}$
@@
Within this approach, the average energy of relativistic neutrinos originating from the source is given by $E_{0}=\sqrt{E_{k}^{2}-m_{k}^{2}}$, where $E_{k}$ denotes the energy and $m_{k}$ the mass of the kk–th mass eigenstate. The quantity $b$, known as the impact parameter, is addressed in Ref. neu18 . As the neutrinos move through the curved background, they approach the central object down to a minimum radial coordinate $r=r_{0}$. In the regime of a weak gravitational field, the equation determining this point of closest approach can be solved analytically using appropriate approximations
@@
$\simeq\frac{{m}_{k}^{2}}{2E_{0}}\left(1+\dfrac{\ell}{2}\right)\Biggl[\sqrt{r_{D}^{2}-r_{0}^{2}}+\sqrt{r_{S}^{2}-r_{0}^{2}}$
@@
$\tilde{\Phi}_{k}$ $\simeq\frac{{m}_{k}^{2}}{2E_{0}}\left(1+\dfrac{\ell}{2}\right)\Biggl[\sqrt{r_{D}^{2}-b^{2}}+\sqrt{r_{S}^{2}-b^{2}}$
@@
$\tilde{\Phi}_{k}=\dfrac{m_{k}^{2}}{2E_{0}}\left(1+\dfrac{\ell}{2}\right)(r_{D}+r_{S})\left(1-\dfrac{b^{2}}{2r_{D}r_{S}}+\dfrac{2M}{r_{D}+r_{S}}\right).$ (40)
As the Lorentz–violating parameter $\ell$ increases, one observes a clear increment in the phase acquired by neutrinos during their journey. For the purposes of this evaluation, the parameters are set to $E_{0}=10\,\mathrm{MeV}$, $r_{D}=10\,\mathrm{km}$, and $r_{S}=10^{5}r_{D}$.
@@
$A_{pq}$ $=\frac{r_{S}+r_{D}}{2E_{0}}\left(1+\dfrac{\ell}{2}\right)\left[1+\dfrac{2M}{r_{D}+r_{S}}-\dfrac{\sum b_{pq}^{2}}{4r_{D}r_{S}}\right],$ (44)
$B_{ij}$ $=-\frac{\sum m_{ij}^{2}}{8E_{0}}\left(1+\dfrac{\ell}{2}\right)\left(\frac{1}{r_{D}}+\frac{1}{r_{S}}\right),$ (45)
@@
[…]
@@ V Gravitational deflection of neutrino trajectories
Now, our analysis turns to the effect of gravitational lensing on neutrino flavor conversion, emphasizing the influence due to $\ell$. Within this lensing configuration, considering a two–flavor neutrino system, the probability of transition from $\nu_{\alpha}$ to $\nu_{\beta}$ is determined under the weak–field approximation, taking into account the spatial arrangement of the emission point, the detector and lensing region neu53 ; neu54 ; Shi:2024flw ; neu65
@@
$\quad\times\left[\exp\left(-\mathbbm{i}\Delta m_{ij}^{2}A_{11}\right)+\exp\left(-\mathbbm{i}\Delta m_{ij}^{2}A_{22}\right)\right]$
@@
$\quad\times\exp\left(-\mathbbm{i}\Delta b_{12}^{2}B_{ij}\right)\exp\left(-\mathbbm{i}\Delta m_{ij}^{2}A_{12}\right)$
@@
$\quad\times\exp\left(\mathbbm{i}\Delta b_{21}^{2}B_{ij}\right)\exp\left(-\mathbbm{i}\Delta m_{ij}^{2}A_{21}\right)\biggr\}.$ (51)
@@
$\quad\times\biggl[\sin^{2}\left(\dfrac{1}{2}\Delta m_{12}^{2}A_{11}\right)+\sin^{2}\left(\dfrac{1}{2}\Delta m_{12}^{2}A_{22}\right)$